If you deposit money today into an account that pays 6.5% interest, how long will it take for you to double your money?
I am having difficulty figuring this out and need some help to solve this problem. Thanks

Answer by Nate(3500)

f(x) = x(1 + r)^t determines amount after t time
f(x) = final amount = 2x because you want your money to double
x = initial amount
r = rate
t = time
2x = x(1 + 0.065)^t
2 = (1.065)^t
log(2) = t*log(1.065)
log(2)/log(1.065) = t
About 11 years (or whatever the time is...)